# AllSpark[Uncaptioned image]: A Multimodal Spatio-Temporal General Intelligence Model with Thirteen Modalities

### v2

@@ V-A Limitations
    Lack of interaction between different modalities. While AllSpark incorporates interactions between infrared and RGB modalities, for other modalities, it essentially uses a unified model for multiple single-modal tasks. This is primarily due to constraints such as the lack of multimodal paired data, the inflexibility of the model architecture, and the limited richness of multimodal interaction tasks. In the future, we plan to explore and work more extensively in these directions.
@@
    There is a lack of in-depth research on multimodal alignment methods. While we introduced the basic principles of the LaRF and aligned all the modalities to the language modality, our approach included only introducing the modality bridge from Lynx as the alignment method. We have not conducted in-depth research into the various methods of multimodal alignment. Moreover, our testing focused on the model’s adaptability to different modalities, and we have not explored potential capabilities such as interactive reasoning, zero-shot learning, or multimodal emergence that the model may possess after alignment with language. In the future, we plan to delve deeper into researching multimodal alignment methods, and we will design additional tasks and experiments to explore and reveal the upper limits and boundaries of large model reasoning capabilities.
@@
    Initial work. Our current work is still in its initial exploratory phase, and we have not carefully refined AllSpark’s adaptability and performance on each modality. As a result, the model exhibits suboptimal performance on certain imaging modalities, such as oblique photography, SAR, and video. In the future, we plan to conduct more refined and targeted adjustments for each modality to enhance overall performance.
@@
    Expensive cost: Multimodal large language models, due to prolonged pretraining, typically possess universal reasoning capabilities in certain modalities. We generalize their applicability to other modalities by utilizing modality bridges to project other modalities onto the language modality. Although we freeze most of the parameters, fine-tuning even once on 2 A6000 GPUs often requires more than a day. Given the increasing training costs for large models, exploring methods to generalize their universal reasoning abilities is one of our future research directions.
  5. 5.
@@ VI Conclusion
Leveraging multimodal data is an inherent requirement for intelligent models to achieve geographic object cognition. Inspired by human cognitive systems and linguistic philosophy, we propose that the construction of multimodal models follow the fundamental principle of Language as Reference Framework (LaRF). Guided by this principle, we use language to balance the cohesion and autonomy of modalities, presenting a unified intelligent model, AllSpark, encompassing thirteen spatio-temporal modalities. The experimental results demonstrated that AllSpark exhibited excellent adaptability and application potential across various spatio-temporal modalities, highlighting the feasibility and potential of constructing multimodal models with LaRF. AllSpark remains an initial exploratory work, and in the future, we aim to delve deeper into the mechanisms guided by natural language, the efficient generalization of large models, and the transition to an interactive paradigm.
TABLE XVI: List of prompts
Dimension
Modal Prompt
1D Text 1.Please determine if this movie review is positive or negative?
@@
2D RGB 1.This remote sensing image belongs to which of the following categories: [Category list of the NWPU dataset] 2.Describe this remote sensing image briefly 3.Find a word that is most relevant to this remote sensing image 4.Describe the key elements in this remote sensing image
MSI 1.Based on the multi-spectral imagery feature description, please classify this object 2.Given the following multi-spectral imagery characteristics, please output the most fitting scene label
HSI 1.Given the spectral information, can you help determine which class this pixel belongs to? 2.Here is the spectral data for a pixel. Considering the typical characteristics of land cover classes, could you provide a detailed analysis and suggest the most likely class for this pixel? 3.The spectral information for a pixel is given, but the data is noisy. Given the potential variability, which land cover classes should be considered as possible candidates for this pixel?
Table 1.Please utilize the provided air quality indicators to accurately predict the concentrations of PM2.5 in the atmosphere 2.Given the air quality indicators, please provide a prediction for the PM2.5 levels at this particular moment
Oblique Photography 1.Please reconstruct the 3D model from the given images of five different views.
Trajectory 1.Based on their past positions and movements in a crowded environment, predict the future trajectory of a selected pedestrian. 2.Using the pedestrian trajectory data, along with additional information about the surrounding environment, predict the future path of the pedestrian. 3.Given the current and past positions of a pedestrian and their neighboring pedestrians, predict the main pedestrian’s trajectory
SAR 1.Based on the SAR imagery feature description, please classify this object 2.Given the following SAR imagery characteristics, please output the most fitting scene label
Infrared 1.Given an infrared image of a person, find and highlight the same person in a set of visible light images, paying close attention to features like clothing, posture, and gait 2.Analyze the characteristics of a person in a visible light image, such as clothing texture, color, and body shape. Then, locate the person with matching features in a series of infrared images
Graph 1.Given the current traffic data including vehicle flow rate, average speed, and time of day from the METRLA dataset, predict the traffic flow 2.Analyze the historical data on vehicle speeds and flow rates from the METRLA dataset for the past week. Identify any patterns or trends and predict the traffic conditions
3D Video 1.Please analyze this video and describe the characteristics of the action in detail 2.Watch this short clip and provide a step-by-step description of the action sequence
Point Cloud 1.Classify the provided point cloud sample into the correct category 2.Look at the point cloud data characteristics and classify the object 3.Please analyze the given point cloud dataset and determine which category it belongs to. Focus on the shape and structure evident in the point cloud

### v3

@@ V-A Limitations
    Lack of interaction between different modalities. AllSpark only facilitates interactions between the language and other modalities, without involving more interactions, such as RGB and point cloud, or hyperspectral and multispectral. This is primarily due to constraints such as the lack of multimodal paired data. However, collecting large-scale paired data for ten modalities is nearly impossible, so we attempt to use language as the alignment reference for each modality, achieving indirect alignment between modalities using unpaired data. AllSpark represents the initial effort in this approach and has demonstrated strong adaptability across various modalities. Also, the adversarial examples also affect the robustness of the proposed model [120]. In the future, we plan to explore and expand our efforts in these directions in the further.
@@
    Initial work. Our current work is still in its initial exploratory phase, and we have not carefully refined AllSpark’s adaptability and performance on each modality. As a result, the model exhibits suboptimal performance on certain modalities, such as oblique photography, and video. The experimental results can be found in Table XIX and Table XVIII. In the future, we plan to conduct more refined and targeted adjustments for each modality to enhance overall performance.
@@
    Expensive cost: Multimodal large language models, due to prolonged pretraining, typically possess universal reasoning capabilities in certain modalities. We generalize their applicability to other modalities by utilizing modality bridges to project other modalities onto the language modality. As shown in Section IV-E, although we freeze most of the parameters, fine-tuning even once on 2 A6000 GPUs often requires more than a day. Given the increasing training costs for large models, exploring methods to generalize their universal reasoning abilities is one of our future research directions.
@@ VI Conclusion
Leveraging multimodal data is an inherent requirement for intelligent models to achieve geographic object cognition. Inspired by human cognitive systems and linguistic philosophy, we propose that the construction of multimodal models follow the fundamental principle of Language as Reference Framework (LaRF). Guided by this principle, we use language to balance the cohesion and autonomy of modalities, presenting a unified intelligent model, AllSpark, encompassing ten spatio-temporal modalities. The experimental results demonstrated that AllSpark exhibited excellent adaptability and application potential across various spatio-temporal modalities, highlighting the feasibility and potential of constructing multimodal models with LaRF. AllSpark remains an initial exploratory work, and in the future, we aim to delve deeper into the mechanisms guided by natural language, the efficient generalization of large models, and the transition to an interactive paradigm.
Appendix A Appendix
A-A List of prompts
The text prompts we used are listed in Table XVII.
TABLE XVII: List of prompts
Dimension Modal Prompts
1D Language 1.Please determine if this movie review is positive or negative?
@@ A-A List of prompts
Table
1.Please utilize the provided air quality indicators to accurately
predict the concentrations of PM2.5 in the atmosphere
2.Given the air quality indicators, please provide a prediction
for the PM2.5 levels at this particular moment
2D RGB
1.This remote sensing image belongs to which of the following
categories: [Category list of the NWPU dataset]
@@
2.Describe this remote sensing image briefly3.Find a word that
is most relevant to this remote sensing image
3.Find a word that is most relevant to this remote sensing image
4.Describe the key elements in this remote sensing image
MSI
1.Based on the multi-spectral imagery feature description, please
classify this object
2.Given the following multi-spectral imagery characteristics, please
output the most fitting scene label
HSI
1.Given the spectral information, can you help determine which
class this pixel belongs to?
2.Here is the spectral data for a pixel. Considering the typical
characteristics of land cover classes, could you provide a detailed
analysis and suggest the most likely class for this pixel?
3.The spectral information for a pixel is given, but the data is
noisy. Given the potential variability, which land cover classes
should be considered as possible candidates for this pixel?
Trajectory
1.Based on their past positions and movements in a crowded
environment, predict the future trajectory of a selected pedestrian
2.Using the pedestrian trajectory data, along with additional
information about the surrounding environment, predict the
future path of the pedestrian
3.Given the current and past positions of a pedestrian and their
neighboring pedestrians, predict the main pedestrian’s trajectory
SAR
1.Based on the SAR imagery feature description, please classify
this object
2.Given the following SAR imagery characteristics, please output
the most fitting scene label
Graph
1.Given the current traffic data including vehicle flow rate, average
speed, and time of day from the METRLA dataset, predict the
traffic flow
2.Analyze the historical data on vehicle speeds and flow rates from
the METRLA dataset for the past week. Identify any patterns or
trends and predict the traffic conditions
3D Point Cloud
1.Classify the provided point cloud sample into the correct category
2.Look at the point cloud data characteristics and classify the object
3.Please analyze the given point cloud dataset and determine which
category it belongs to. Focus on the shape and structure evident in
the point cloud
A-B Video
For the video modality, we evaluated AllSpark’s performance on action recognition tasks using the UCF101[121] dataset. The UCF101 dataset is a human action recognition dataset comprising 101 action classes with a total of 13,320 video clips. The videos have a combined duration of 27 hours and a resolution of 320*240 pixels and were sourced from YouTube.
The model is tasked with understanding videos and accurately classifying them into one of the 101 classes, with the evaluation metric being the top-1 accuracy. In Table XVIII, we compare AllSpark with the current state-of-the-art models.
Currently, AllSpark’s adaptation to the video modality is not optimal, as it shows a significant difference from the baseline model results. We attribute this to two main reasons: 1. The high redundancy in video information increases the training cost for AllSpark. We trained it for only 3 epochs on the dataset. 2. AllSpark’s model architecture lacks flexibility for three-dimensional data, making it less effective at capturing temporal information.
TABLE XVIII: video classification with allspark
Method Acc(%)
OPN 59.6
VCOP 72.4
SpeedNet 81.1
VTHCL 82.1
CVRL 94.4
VideoMAE v1 96.1
VideoMAE v2 99.6
AllSpark 27.5
A-C Oblique Photography
For the oblique photography modality, we tested AllSpark’s performance on the 3D reconstruction task using the WHU-OMVS[122] dataset. WHU-OMVS is an oblique photography dataset designed for 3D reconstruction tasks. The dataset provides imagery from five different viewpoints, along with camera parameters and other relevant information. It consists of six areas, and in our experiments, area 1 is used as the training set, and area 2 is used as the test set.
The model takes five-view images as input, and the goal is to output depth maps for reconstructing 3D models. The evaluation metric used was the percentage of accurate grids in total (PAG)[122], calculated by the following formula:
$PAG_{a}=\left(\frac{m_{a}}{m}\right)*100$ (7)
The suffixes in the PAG represent different accuracy standards, where ${PAG}_{6}$ signifies an error within 0.6 metres and ${PAG}_{10}$ indicates an error within 1 metre. Table XIX compares AllSpark with popular multiview 3D reconstruction models on the WHU-OMVS dataset.
AllSpark falls short in terms of accuracy compared to modality-specific expert models. We speculate two possible reasons: 1. Dense spatial information gradually diminishes in the deep structure of large models, a point verified in our exploratory experiments involving segmentation, detection, etc.; 2. The model architecture lacks flexibility and cannot connect gradual features like expert 3D reconstruction models such as Ada-MVS. Additionally, these methods lack the ability to design specific model structures for processing, severely restricting their performance. Adapting multimodal large models to dense prediction tasks and optimizing the architecture are future research directions.
TABLE XIX: 3D reconstruction with allspark
Method ${PAG}_{6}$ ${PAG}_{10}$
MVSNet 81.15 91.44
CasMVSNet 95.45 98.02
Ada-MVS 96.14 98.1
UCSNet 96.25 98.45
AllSpark 6.4 10.4